OpenStudy (anonymous):

write the equation of a line that includes the point (4,-2) and has a slope of 0 in slope intercept form.

OpenStudy (anonymous):

y=-2

OpenStudy (yrelhan4):

y-y1/x-x1 = m.. where x1,y1 are the points and m is the slope. just substitute the values.

OpenStudy (anonymous):

slope intercept form is y=mx+b where m is the slope and b is the y-intercept

OpenStudy (anonymous):

y=0(4)+(-2)

OpenStudy (anonymous):

y=0+(-2) y=-2